OpenStudy (anonymous):

Limit as x->(pi/2) of [tanx + 1/(x-pi/2)]? Anyone?

OpenStudy (anonymous):

do you get zero over zero?

OpenStudy (anonymous):

i think you are supposed to recognize this as the derivative of tangent, evaluated at \(\frac{\pi}{4}\)

OpenStudy (anonymous):

oops rather \(\frac{\pi}{2}\)

OpenStudy (anonymous):

derivative of tangent is secant squared, and secant square of \(\frac{\pi}{2}\) is your answer